xx IOHDF5Util

IOHDF5Util

Thomas Radke

Date

Abstract

Thorn IOHDF5Util provides functionality to implement HDF5 I/O in Cactus.

1 Purpose

Thorn IOHDF5Util provides common routines which can be used to implement Cactus I/O methods for input/output of Cactus variables in the HDF5 standard (Hierarchical Data Format Version 5).

Currently, IOHDF5Util is inherited by the IOHDF5 and IOStreamHDF5 I/O thorns which share its routines to implement their own I/O methods doing parallel HDF5 file I/O and online HDF5 data streaming.

2 Building A Cactus Configuration with IOHDF5Util The Cactus distribution does not contain the HDF5 header files and library which is used by thorn IOHDF5Util. So you need to configure it as an external software package via:

xxmakex<configuration>-configxHDF5=yes
xxxxxxxxxxxxxxxxxxxxxxxxxxxxx[HDF5_DIR=<pathxtoxHDF5xpackage>]

The configuration script will look in some default places for an installed HDF5 package. If nothing is found this way you can explicitly specify it with the HDF5xDIR configure variable.

3 Parameters

11compressionxlevel 10Scope: privatex01  INT

11Description: Compression level to use for writing HDF5 data

11 Range 10 x01 Default: (none)
110:9  10Higher numbers compress better, a value of zero disables compression

11

11outxfileinfo  10Scope: shared from IOx01KEYWORD

11

11verbose  10Scope: shared from IOx01KEYWORD

11

4 Interfaces

General Implements: iohdf5util

Inherits: io 5 Schedule This section lists all the variables which are assigned storage by thorn CactusPUGHIO/IOHDF5Util. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE

Scheduled Functions

CCTKxSTARTUP x iohdf5utilxstartup x iohdf5util startup routine x

11x11After: 11ioutilxstartup
11 x 11Language:11c
11 x 11Type: 11function

11

CCTKxTERMINATE x iohdf5utilxterminate x iohdf5util termination routine x

11x11Before: 11driverxterminate
11 x 11Language:11c
11 x 11Type: 11function